OpenStudy (adamconner):

what are the zeros for 6(x-1) ?

OpenStudy (adamconner):

is it just x=1?

OpenStudy (whpalmer4):

Yes! The zeros of an expression are the values that make the expression equal 0. \[6(x-1) = 0\]Divide both sides by 6\]\[(x-1) = 0\]\[x-1+1 = 0+1\]\[x=1\]